A ladder makes alpha angle with the ground. Its bottom is pulled by a And due to this it slips by b with the wall and makes beta angle with the wall then a is equal to

a. b tan(alpha+beta)/2

b. b tan(alpha+beta)

c. b tan( ( alpha+beta)/2)

d. (b tan((alpha+beta)/2))/2


1
Expert's answer
2021-02-24T07:49:57-0500

Worth noting, the data in problem don't correspond to the answers. In answers the two angles are symmetric, so "\alpha" and "\beta" should be the angles with the ground.

Let l be the length of the ladder. Therefore, the distance of the bottom from the wall is "l\cos\alpha" and the height of top is "l\sin\alpha." If we pull the bottom by a, the distance from the wall will be "l\cos\beta= l\cos\alpha + a" and the height of the top will be "l\sin\beta =l\sin\alpha - b."

Therefore,

"l\cos\beta= l\cos\alpha + a, \\
l\sin\beta =l\sin\alpha - b."

So "l = \dfrac{b}{\sin\alpha - \sin\beta}" and "a =b\cdot \dfrac{\cos\beta-\cos\alpha}{\sin\alpha - \sin\beta} = b\cdot\dfrac{-2\sin\frac{\alpha+\beta}{2}\sin\frac{-\alpha+\beta}{2}}{2\sin\frac{\alpha-\beta}{2}\cos\frac{\alpha+\beta}{2}} =b \tan\dfrac{\alpha+\beta}{2}" .